# Electricity Demand and Grid Impacts of AI Data Centers: Challenges and Prospects

### v3

@@ IV-B1 Impacts on Power Dispatch and Reserve Scheduling
Maintaining the balance between generation and demand is a fundamental objective of power system operation. On the short‑term timescale, power balance is achieved through unit commitment (UC) and economic dispatch (ED), supported by reserve scheduling to accommodate unforeseen power variations [82]. The immense and bursty power usage of AI data centers, particularly during intensive model training and large-scale inference, introduces significant uncertainty and rapid load variations to the grid operation [17]. This challenges conventional deterministic power dispatch schemes and necessitates advanced short‑term forecasting techniques for AI data center loads [17, 83]. However, accurate prediction remains inherently difficult due to their complex, non‑cyclic patterns and dynamic scheduling behaviors of AI workloads. In addition, the rapid and unpredictable ramping of AI data center loads imposes greater reserve requirements. To ensure reliability amid large and fast load variations, grid operators need to maintain elevated spinning and non‑spinning reserve margins, which increase operational costs and call for additional fast‑responding reserve resources.
@@ IV-B2 Impacts on Electricity Market and Price
Large-scale AI data center loads are expected to place upward pressure on capacity market clearing prices, as higher peak demands require additional capacity commitments. This impact has been observed in PJM, where capacity market clearing prices for the 2026-2027 delivery year increased to \$329.17/MW, over ten times higher than the price of \$28.92/MW in the 2024-2025 delivery year, with rapid data center growth identified as a major contributing factor [84, 85]. Similar effects are emerging in New York ISO and ERCOT, where anticipated data center expansion is influencing forward capacity markets and resource adequacy planning [86, 87]. In addition, the substantial and often unpredictable demand from AI data centers raises marginal generation costs in wholesale markets, as more expensive mid-merit and peaking units are dispatched to meet AI-driven loads [14]. It also amplifies price volatility in day-ahead and real-time markets due to short-term large AI load fluctuations [88]. Rapid AI load surges can result in sharp increases in locational marginal prices (LMPs), particularly in transmission-congested areas [89].
@@ IV-C1 Grid Disturbances Ride-Through
AI data centers interface with the power grid through high-capacity power electronic converters, which are sensitive to short-duration voltage and frequency disturbances at the point of interconnection, such as voltage sags, voltage dips, and upward or downward frequency deviations. These disturbances can induce substantial power oscillations and even trigger the disconnection of an AI data center from the grid to protect internal electronic devices. Without adequate fault ride-through (FRT) capabilities or grid disturbance-tolerance measures, such sudden trips can occur during grid faults, potentially exacerbating the original grid disturbance. In regions with high concentrations of AI data centers, simultaneous tripping events can take place and lead to cascading outages [90]. These impacts have been observed in practice. For example, high-frequency undamped power oscillations appeared in a data center-rich region in Dominion Energy’s grid, triggered by a once-a-second voltage sag [91]. ERCOT has also reported multiple events wherein large computing loads tripped during transmission faults due to insufficient ride-through capability [92]. For example, Figure 5 illustrates the grid frequency dynamics under several large electronic load trip events studied in ERCOT [93]. It shows that the loss of more than 2600 MW of such loads can drive system frequency up to 60.4 Hz, a level at which conventional generators may fail to ride through, potentially leading to uncontrolled cascading events [93]. To this end, ERCOT considered imposing ride-through requirements for large data center and crypto-mining loads [94] to ensure reliable interconnection and grid operation.
Figure 5: ERCOT system frequency response to several large electronic load trip events [93].
@@ IV-C2 Power Fluctuations and Grid Stability
AI data centers can introduce rapid power fluctuations during cold starts, planned shutdowns, workload transitions, or sudden load interruptions triggered by faults or operational contingencies. For example, training large-scale AI models across tens of thousands of GPUs can induce significant power swings, arising from the synchronous nature of these jobs and the alternation between a power-intensive computation phase and a less power-demanding communication phase [18]. In addition, inference workloads often exhibit rapid “start-stop” behavior with sharp spikes followed by sudden drops in power consumption, as shown in Figure 3. Unlike conventional electromechanical loads, power electronics-based AI data center demand may change by tens to hundreds of megawatts within sub-second intervals, posing substantial stability risks to the power grid. Such abrupt power variations can challenge the grid’s frequency stability [95, 96], induce voltage excursions [97], and interact with the dynamics of grid-connected generation and control systems, potentially exciting oscillations and instability [98]. For instance, a major resonance event was reported in 2017 across multiple Facebook data centers and analyzed in [99]. As noted in [100], unlike solar plants and wind farms, data centers typically have internal impedances much higher than those of transmission lines or distribution transformers, which can amplify such resonance phenomena. From a data center perspective, reference [18] proposes three classes of power stabilization techniques to mitigate large power swings, including (i) software-based methods that inject controlled workloads to smooth power transitions, (ii) GPU-level firmware mechanisms that enforce ramping constraints and power floors, and (iii) rack-level energy storage systems to reduce fluctuations by absorbing and releasing power as needed. Beyond these approaches, addressing the challenge of large power fluctuations also requires accurate dynamic modeling of AI data center behavior and close coordination between grid operators and data center operators to ensure effective mitigation and maintain system-level stability. Moreover, as data centers typically employ fast-acting UPS systems for power continuity, transitions between grid supply and UPS operation during disturbances must be carefully managed to prevent large transient power swings or synchronization issues upon reconnection.
@@ IV-C3 Power Quality Issues
High‑frequency switching devices and power electronic equipment in data centers can introduce significant harmonic distortions to the grid, degrading local power quality and posing risks to nearby electrical facilities [101]. These distortions originate primarily from non‑linear power supplies, AC/DC conversions, and the control of server PFC circuits [102]. In large-scale AI facilities, where thousands of server racks operate simultaneously, these harmonics can accumulate and propagate into the local distribution system, presenting risks to industrial and residential equipment, such as malfunctions, accelerated wear, overheating, and electrical fires. According to a 2024 Bloomberg analysis [21], sensor data from more than 700,000 homes reveal a strong correlation between proximity to data centers and declining power quality, with evident effects observed within 20 miles of major data center clusters. Beyond harmonic distortion, AI data centers may contribute to other power quality issues, such as voltage flicker during rapid load changes, unbalanced loads across three phases [97], inter-harmonics caused by certain switching patterns in high‑frequency converters [103]. These power quality issues can propagate through the grid, potentially degrading protection system performance, reducing metering accuracy, shortening equipment lifespans, and impairing control system stability [97]. To mitigate power quality issues and maintain compliance with industry standards such as IEEE Standard 519‑2014 [102], key approaches include: (i) active harmonic filtering at the AI data center facilities, (ii) phase balancing and load redistribution to minimize neutral current distortion, (iii) coordination with grid operators for harmonic monitoring and mitigation planning, and (iv) UPS and converter design optimization to minimize harmonic generation at the source. As AI data center capacity keeps expanding, proactive power quality management will be critical for maintaining grid stability and ensuring equipment safety and regulatory compliance.
@@ IV-D1 Cybersecurity
AI data centers not only present major physical impacts on the power grid, as discussed above, but also introduce emerging cybersecurity risks to power systems. Their highly concentrated electricity demand, dependence on programmable power electronics, and integration with cloud-based control architectures collectively increase the vulnerability of power systems to cyber attacks. Specifically, coordinated cyber intrusions could manipulate workload scheduling or UPS switching, deliberately inducing sudden surges or drops of hundreds of megawatts in power demand. Such abrupt load changes pose a significant threat to grid stability, potentially overloading local distribution networks, disrupting frequency regulation, and triggering cascading failures across interconnected systems [104]. For instance, in 2024, a large-scale disconnection event occurred in Virginia’s “Data Center Alley”, where a protection system failure caused 60 out of more than 200 data centers to suddenly disconnect from the grid and transition to on-site generators [105]. This abrupt loss of load introduced a significant power imbalance, compelling the utility to rapidly curtail generation in order to avert cascading outages. Although the event was not triggered by malicious cyber attacks, it underscores how sudden shifts in data center loads can impose substantial stress on grid stability.
Beyond transient disturbances, maliciously orchestrated or malfunction-induced workload patterns can act as sources of forced oscillations, periodically modulating AI data center demand in ways that resonate with the grid’s natural modes. Such oscillatory injections can amplify inter-area oscillations, reduce damping margins, and in extreme cases, trigger widespread instability [106]. Unlike conventional industrial loads, the inherent controllability and programmability of AI workloads render them particularly susceptible to being exploited as virtual power perturbation sources. This tight coupling between cyber-level control and physical load response highlights the urgent need for integrated approaches that jointly address cybersecurity and power system dynamics. Effective countermeasures include the implementation of secure communication protocols, the development of anomaly detection mechanisms tailored to workload-induced oscillations, and the establishment of collaborative frameworks between data center operators and grid utilities to prevent cascading failures triggered via cyber attacks.
@@ IV-D2 Decarbonization
According to the IEA report [8], carbon emissions from the electricity use of global data centers reached 180 million tonnes (Mt) in 2024 and are projected to rise to 300 Mt by 2035, raising concerns that the rapid growth of AI facilities could further exacerbate climate change [107]. Nevertheless, AI data centers generally do not produce direct carbon emissions during normal operation, apart from on-site backup generators such as diesel units, which are only infrequently activated. Instead, AI data centers mainly account for the indirect carbon footprint [^2] Accordingly, the Greenhouse Gas (GHG) Protocol [108, 109] defines two categories of emissions, Scope 1 and Scope 2, to distinguish between direct and indirect (or attributed) emissions for carbon accounting. associated with their electricity consumption, with the carbon emission intensities determined by the generation fuel mix and electric power flow of the supplying power grid [110]; see [111] for a more detailed introduction. Consequently, the coupling between electricity demand and grid carbon intensity highlights the importance of considering both energy efficiency measures within AI data centers and the decarbonization trajectory of power grids [112]. At present, data center owners achieve their decarbonization goals primarily by co-locating renewable generation with data center campuses or through procuring renewable energy in the electricity market via power purchase agreements (PPAs) [113] or renewable energy certificates (RECs) [114]. For example, Google launched its “24/7 Carbon-Free Energy” program [115] aimed at sourcing carbon-free electricity for its data center operations on a continuous 24 hours-7 days time basis for decarbonization.
@@ IV-D3 Water Resources
Many large-scale data centers utilize evaporative cooling systems to dissipate heat, a process that consumes substantial volumes of water daily, often amounting to tens of thousands or even millions of liters of water depending on the facility size and climate conditions. As indicated in report [8], a 100 MW data center in the U.S. on average consumes around 2 million liters of water per day, equivalent to about 6,500 households. It estimates that global water consumption for data centers is currently around 560 billion liters per year and could rise to around 1,200 billion liters per year in 2030. As a consequence, together with high electricity demand, substantial water withdrawals are emerging as a critical limiting factor for AI-driven data center expansion, especially during summer peaks when both water availability and power supply are under stress. According to Bloomberg [116], nearly two-thirds of new U.S. data centers built since 2022 are located in high water-stress regions, such as the Southwest, Texas, and parts of the Midwest, some of which face growing risks of drought and water scarcity. Reference [117] proposes a general framework for evaluating the water impact of computing by factoring in both spatial and temporal variations in water stress. In [118], the factors influencing data center water use are analyzed, such as server efficiency, server utilization, and cooling techniques. The study also finds that there is no single recipe for minimizing water use; instead, optimal outcomes depend on tailored combinations of these factors. To reduce stress on water resources, integrated planning frameworks that jointly assess electricity and water footprints of AI data centers are necessary. Effective strategies include adopting low-water cooling technologies in water-scarce regions, siting facilities with attention to both water availability and grid capacity, and incorporating climate resilience metrics into future AI infrastructure planning.
@@ V-A1 Improving Forecasting of AI Load
Accurate forecasting of AI data center loads, particularly during large-scale model training and inference stages, is critical for grid planning and operation. Unlike traditional data center loads, AI workloads exhibit highly irregular temporal patterns, with abrupt transitions between low-utilization and peak-demand periods. These rapid changes can generate sharp power surges that challenge power dispatch, reserve scheduling, and real-time grid stability. Effective forecasting requires a comprehensive understanding of AI workload characteristics, including job scheduling policies, training cycle durations, hardware utilization patterns, and the interaction between compute intensity and cooling system demand [119]. To address privacy concerns, load forecasting can be performed internally by AI data centers, while only aggregated results are reported to the grid operator. Alternatively, collaborative forecasting frameworks, where AI data center operators share anonymized workload information with grid operators, can be developed to enhance situational awareness and support preventive grid operations that mitigate potential impacts. As a complement, recent advances in machine learning (ML) techniques can be leveraged to enhance AI load forecasting [120, 17]. These approaches integrate telemetry data from IT infrastructure, job queue statistics, and environmental factors with advanced ML tools to predict short-term fluctuations and identify potential peak demand events before they occur. For example, hybrid statistical-ML models have been shown to improve short-term prediction accuracy for cloud and AI computing facilities compared to conventional auto-regressive models [83].

### v4

@@ IV-B1 Impacts on Power Dispatch and Reserve Scheduling
Maintaining the balance between generation and demand is a fundamental objective of power system operation. On the short‑term timescale, power balance is achieved through unit commitment (UC) and economic dispatch (ED), supported by reserve scheduling to accommodate unforeseen power variations [wood2013power]. The immense and bursty power usage of AI data centers, particularly during intensive model training and large-scale inference, introduces significant uncertainty and rapid load variations to the grid operation [mughees2025forecast]. This challenges conventional deterministic power dispatch schemes and necessitates advanced short‑term forecasting techniques for AI data center loads [mughees2025forecast, li2023shortterm]. However, accurate prediction remains inherently difficult due to their complex, non‑cyclic patterns and dynamic scheduling behaviors of AI workloads. In addition, the rapid and unpredictable ramping of AI data center loads imposes greater reserve requirements. To ensure reliability amid large and fast load variations, grid operators need to maintain elevated spinning and non‑spinning reserve margins, which increase operational costs and call for additional fast‑responding reserve resources.
@@ IV-B2 Impacts on Electricity Market and Price
Large-scale AI data center loads are expected to place upward pressure on capacity market clearing prices, as higher peak demands require additional capacity commitments. This impact has been observed in PJM, where capacity market clearing prices for the 2026-2027 delivery year increased to \$329.17/MW, over ten times higher than the price of \$28.92/MW in the 2024-2025 delivery year, with rapid data center growth identified as a major contributing factor [pjm2025capacity, kunkel2025pjm]. Similar effects are emerging in New York ISO and ERCOT, where anticipated data center expansion is influencing forward capacity markets and resource adequacy planning [nyiso2025powertrends, ercot2024longterm]. In addition, the substantial and often unpredictable demand from AI data centers raises marginal generation costs in wholesale markets, as more expensive mid-merit and peaking units are dispatched to meet AI-driven loads [aljbour2024powering]. It also amplifies price volatility in day-ahead and real-time markets due to short-term large AI load fluctuations [businessinsider2025]. Rapid AI load surges can result in sharp increases in locational marginal prices (LMPs), particularly in transmission-congested areas [JLARC2024DataCenter].
@@ IV-C1 Grid Disturbances Ride-Through
AI data centers interface with the power grid through high-capacity power electronic converters, which are sensitive to short-duration voltage and frequency disturbances at the point of interconnection, such as voltage sags, voltage dips, and upward or downward frequency deviations. These disturbances can induce substantial power oscillations and even trigger the disconnection of an AI data center from the grid to protect internal electronic devices. Without adequate fault ride-through (FRT) capabilities or grid disturbance-tolerance measures, such sudden trips can occur during grid faults, potentially exacerbating the original grid disturbance. In regions with high concentrations of AI data centers, simultaneous tripping events can take place and lead to cascading outages [ieee_frt_2018]. These impacts have been observed in practice. For example, high-frequency undamped power oscillations appeared in a data center-rich region in Dominion Energy’s grid, triggered by a once-a-second voltage sag [mishra2025understanding]. ERCOT has also reported multiple events wherein large computing loads tripped during transmission faults due to insufficient ride-through capability [Gravois2024_ERCO]. For example, Figure 5 illustrates the grid frequency dynamics under several large electronic load trip events studied in ERCOT [Jeff2025_ERCO]. It shows that the loss of more than 2600 MW of such loads can drive system frequency up to 60.4 Hz, a level at which conventional generators may fail to ride through, potentially leading to uncontrolled cascading events [Jeff2025_ERCO]. To this end, ERCOT considered imposing ride-through requirements for large data center and crypto-mining loads [ERCOT2025_MB062325] to ensure reliable interconnection and grid operation.
Figure 5: ERCOT system frequency response to several large electronic load trip events [Jeff2025_ERCO].
@@ IV-C2 Power Fluctuations and Grid Stability
AI data centers can introduce rapid power fluctuations during cold starts, planned shutdowns, workload transitions, or sudden load interruptions triggered by faults or operational contingencies. For example, training large-scale AI models across tens of thousands of GPUs can induce significant power swings, arising from the synchronous nature of these jobs and the alternation between a power-intensive computation phase and a less power-demanding communication phase [choukse2025power]. In addition, inference workloads often exhibit rapid “start-stop” behavior with sharp spikes followed by sudden drops in power consumption, as shown in Figure 3. Unlike conventional electromechanical loads, power electronics-based AI data center demand may change by tens to hundreds of megawatts within sub-second intervals, posing substantial stability risks to the power grid. Such abrupt power variations can challenge the grid’s frequency stability [kundur_power_system_stability_1994, ulbig2014impact], induce voltage excursions [bollen_power_quality_2000], and interact with the dynamics of grid-connected generation and control systems, potentially exciting oscillations and instability [meegahapola2020power]. For instance, a major resonance event was reported in 2017 across multiple Facebook data centers and analyzed in [LowFrequencyModel]. As noted in [ImpedanceModel], unlike solar plants and wind farms, data centers typically have internal impedances much higher than those of transmission lines or distribution transformers, which can amplify such resonance phenomena. From a data center perspective, reference [choukse2025power] proposes three classes of power stabilization techniques to mitigate large power swings, including (i) software-based methods that inject controlled workloads to smooth power transitions, (ii) GPU-level firmware mechanisms that enforce ramping constraints and power floors, and (iii) rack-level energy storage systems to reduce fluctuations by absorbing and releasing power as needed. Beyond these approaches, addressing the challenge of large power fluctuations also requires accurate dynamic modeling of AI data center behavior and close coordination between grid operators and data center operators to ensure effective mitigation and maintain system-level stability. Moreover, as data centers typically employ fast-acting UPS systems for power continuity, transitions between grid supply and UPS operation during disturbances must be carefully managed to prevent large transient power swings or synchronization issues upon reconnection.
@@ IV-C3 Power Quality Issues
High‑frequency switching devices and power electronic equipment in data centers can introduce significant harmonic distortions to the grid, degrading local power quality and posing risks to nearby electrical facilities [3dfs2015harmonics]. These distortions originate primarily from non‑linear power supplies, AC/DC conversions, and the control of server PFC circuits [ieee519]. In large-scale AI facilities, where thousands of server racks operate simultaneously, these harmonics can accumulate and propagate into the local distribution system, presenting risks to industrial and residential equipment, such as malfunctions, accelerated wear, overheating, and electrical fires. According to a 2024 Bloomberg analysis [bloomberg2024ai], sensor data from more than 700,000 homes reveal a strong correlation between proximity to data centers and declining power quality, with evident effects observed within 20 miles of major data center clusters. Beyond harmonic distortion, AI data centers may contribute to other power quality issues, such as voltage flicker during rapid load changes, unbalanced loads across three phases [bollen_power_quality_2000], inter-harmonics caused by certain switching patterns in high‑frequency converters [arrillaga_power_quality_2000]. These power quality issues can propagate through the grid, potentially degrading protection system performance, reducing metering accuracy, shortening equipment lifespans, and impairing control system stability [bollen_power_quality_2000]. To mitigate power quality issues and maintain compliance with industry standards such as IEEE Standard 519‑2014 [ieee519], key approaches include: (i) active harmonic filtering at the AI data center facilities, (ii) phase balancing and load redistribution to minimize neutral current distortion, (iii) coordination with grid operators for harmonic monitoring and mitigation planning, and (iv) UPS and converter design optimization to minimize harmonic generation at the source. As AI data center capacity keeps expanding, proactive power quality management will be critical for maintaining grid stability and ensuring equipment safety and regulatory compliance.
@@ IV-D1 Cybersecurity
AI data centers not only present major physical impacts on the power grid, as discussed above, but also introduce emerging cybersecurity risks to power systems. Their highly concentrated electricity demand, dependence on programmable power electronics, and integration with cloud-based control architectures collectively increase the vulnerability of power systems to cyber attacks. Specifically, coordinated cyber intrusions could manipulate workload scheduling or UPS switching, deliberately inducing sudden surges or drops of hundreds of megawatts in power demand. Such abrupt load changes pose a significant threat to grid stability, potentially overloading local distribution networks, disrupting frequency regulation, and triggering cascading failures across interconnected systems [maleki2025survey]. For instance, in 2024, a large-scale disconnection event occurred in Virginia’s “Data Center Alley”, where a protection system failure caused 60 out of more than 200 data centers to suddenly disconnect from the grid and transition to on-site generators [McLaughlin_2025]. This abrupt loss of load introduced a significant power imbalance, compelling the utility to rapidly curtail generation in order to avert cascading outages. Although the event was not triggered by malicious cyber attacks, it underscores how sudden shifts in data center loads can impose substantial stress on grid stability.
Beyond transient disturbances, maliciously orchestrated or malfunction-induced workload patterns can act as sources of forced oscillations, periodically modulating AI data center demand in ways that resonate with the grid’s natural modes. Such oscillatory injections can amplify inter-area oscillations, reduce damping margins, and in extreme cases, trigger widespread instability [sun2018cyber]. Unlike conventional industrial loads, the inherent controllability and programmability of AI workloads render them particularly susceptible to being exploited as virtual power perturbation sources. This tight coupling between cyber-level control and physical load response highlights the urgent need for integrated approaches that jointly address cybersecurity and power system dynamics. Effective countermeasures include the implementation of secure communication protocols, the development of anomaly detection mechanisms tailored to workload-induced oscillations, and the establishment of collaborative frameworks between data center operators and grid utilities to prevent cascading failures triggered via cyber attacks.
@@ IV-D2 Decarbonization
According to the IEA report [iea2024energyai], carbon emissions from the electricity use of global data centers reached 180 million tonnes (Mt) in 2024 and are projected to rise to 300 Mt by 2035, raising concerns that the rapid growth of AI facilities could further exacerbate climate change [Gibney2022Jul]. Nevertheless, AI data centers generally do not produce direct carbon emissions during normal operation, apart from on-site backup generators such as diesel units, which are only infrequently activated. Instead, AI data centers mainly account for the indirect carbon footprint [^2] Accordingly, the Greenhouse Gas (GHG) Protocol [world2004greenhouse, world2014scope2] defines two categories of emissions, Scope 1 and Scope 2, to distinguish between direct and indirect (or attributed) emissions for carbon accounting. associated with their electricity consumption, with the carbon emission intensities determined by the generation fuel mix and electric power flow of the supplying power grid [chen2024carbon]; see [chen2024towards] for a more detailed introduction. Consequently, the coupling between electricity demand and grid carbon intensity highlights the importance of considering both energy efficiency measures within AI data centers and the decarbonization trajectory of power grids [chen2024enhance]. At present, data center owners achieve their decarbonization goals primarily by co-locating renewable generation with data center campuses or through procuring renewable energy in the electricity market via power purchase agreements (PPAs) [kansal2018introduction] or renewable energy certificates (RECs) [lau2008bottom]. For example, Google launched its “24/7 Carbon-Free Energy” program [google247program] aimed at sourcing carbon-free electricity for its data center operations on a continuous 24 hours-7 days time basis for decarbonization.
@@ IV-D3 Water Resources
Many large-scale data centers utilize evaporative cooling systems to dissipate heat, a process that consumes substantial volumes of water daily, often amounting to tens of thousands or even millions of liters of water depending on the facility size and climate conditions. As indicated in report [iea2024energyai], a 100 MW data center in the U.S. on average consumes around 2 million liters of water per day, equivalent to about 6,500 households. It estimates that global water consumption for data centers is currently around 560 billion liters per year and could rise to around 1,200 billion liters per year in 2030. As a consequence, together with high electricity demand, substantial water withdrawals are emerging as a critical limiting factor for AI-driven data center expansion, especially during summer peaks when both water availability and power supply are under stress. According to Bloomberg [Nicoletti2025AIWater], nearly two-thirds of new U.S. data centers built since 2022 are located in high water-stress regions, such as the Southwest, Texas, and parts of the Midwest, some of which face growing risks of drought and water scarcity. Reference [wu2025not] proposes a general framework for evaluating the water impact of computing by factoring in both spatial and temporal variations in water stress. In [Lei2025Jun], the factors influencing data center water use are analyzed, such as server efficiency, server utilization, and cooling techniques. The study also finds that there is no single recipe for minimizing water use; instead, optimal outcomes depend on tailored combinations of these factors. To reduce stress on water resources, integrated planning frameworks that jointly assess electricity and water footprints of AI data centers are necessary. Effective strategies include adopting low-water cooling technologies in water-scarce regions, siting facilities with attention to both water availability and grid capacity, and incorporating climate resilience metrics into future AI infrastructure planning.
@@ V-A1 Improving Forecasting of AI Load
Accurate forecasting of AI data center loads, particularly during large-scale model training and inference stages, is critical for grid planning and operation. Unlike traditional data center loads, AI workloads exhibit highly irregular temporal patterns, with abrupt transitions between low-utilization and peak-demand periods. These rapid changes can generate sharp power surges that challenge power dispatch, reserve scheduling, and real-time grid stability. Effective forecasting requires a comprehensive understanding of AI workload characteristics, including job scheduling policies, training cycle durations, hardware utilization patterns, and the interaction between compute intensity and cooling system demand [cupelli2018data]. To address privacy concerns, load forecasting can be performed internally by AI data centers, while only aggregated results are reported to the grid operator. Alternatively, collaborative forecasting frameworks, where AI data center operators share anonymized workload information with grid operators, can be developed to enhance situational awareness and support preventive grid operations that mitigate potential impacts. As a complement, recent advances in machine learning (ML) techniques can be leveraged to enhance AI load forecasting [Ma2025May, mughees2025forecast]. These approaches integrate telemetry data from IT infrastructure, job queue statistics, and environmental factors with advanced ML tools to predict short-term fluctuations and identify potential peak demand events before they occur. For example, hybrid statistical-ML models have been shown to improve short-term prediction accuracy for cloud and AI computing facilities compared to conventional auto-regressive models [li2023shortterm].